The angles of a pentagon are x, 2x, x+30, Xtake away 10, x+40degrees find the value of x
Answer by stanbon(75887)

The angles of a pentagon are x, 2x, x+30, Xtake away 10, x+40degrees find the value of x
---------------
Sum of angles of a pentagon = (5-2)180 = 3*180 = 540 degrees
-----
Equation:
x + 2x + x+30 + x-10 + x+40 = 540
6x + 60 = 540
x + 10 = 90
-----------------
x = 80 degrees
Cheers,
Stan H.
-----------